# Large Language Models as Span Annotators

### v2

@@ Abstract
Span annotation is the task of localizing and classifying text spans according to custom guidelines. Annotated spans can be used to analyze and evaluate high-quality texts for which single-score metrics fail to provide actionable feedback. Until recently, span annotation was limited to human annotators or fine-tuned models. In this study, we show that large language models (LLMs) can serve as flexible and cost-effective span annotation backbones. To demonstrate their utility, we compare LLMs to skilled human annotators on three diverse span annotation tasks: evaluating data-to-text generation, identifying translation errors, and detecting propaganda techniques. We demonstrate that LLMs achieve inter-annotator agreement (IAA) comparable to human annotators at a fraction of a cost per output annotation. We also manually analyze model outputs, finding that LLMs make errors at a similar rate to human annotators. We release the dataset of more than 40k model and human annotations for further research. [^1] Project website: https://llm-span-annotators.github.io
@@ 1 Introduction
Fine-grained aspects of texts, such as faithfulness or coherence, depend on local lexical choices. In order to reflect these aspects in quality judgments of texts, we need analysis techniques that provide the appropriate amount of detail. However, most automatic evaluation metrics for natural language generation (NLG) are based on direct assessment, producing holistic scores for each evaluated aspect Gkatzia & Mahamood (2015); Sai et al. (2023); Schmidtová et al. (2024). Although numerical values make it easy to rank systems, these metrics are too simplistic and susceptible to biases or miscalibration of the underlying models Gehrmann et al. (2023); Gao et al. (2024); Liu et al. (2024); Wang et al. (2024).
The subject of our study, span annotation (Figure 1), offers an alternative approach. Instead of assigning a single score for each evaluated aspect, the goal of span annotation is to localize text spans and classify them according to task-specific guidelines. In contrast to numerical rating, annotations are aligned to the evaluated text, which makes them more explainable and actionable. The annotations can also be examined post hoc, which brings more accountability into the evaluation process.
Despite its advantages, span annotation has not yet been widely applied in automatic NLG evaluation. The method traditionally required human annotators, making it costly and difficult to scale Da San Martino et al. (2019); Thomson & Reiter (2020); Popovic (2020); Kocmi et al. (2024c). The situation may change with the upcoming LLM-as-a-judge paradigm, in which LLMs follow instructions for the evaluation process similarly to human annotators. However, current LLM-as-a-judge applications still focus mainly on document-level classification or numerical ratings on the Likert scale Zheng et al. (2023); Gu et al. (2024).
In this work, we investigate whether LLMs can be applied as a substitute for human annotators on span annotation tasks. First, we focus on how to produce span annotations with LLMs. We discuss input tagging methods and experiment with prompting techniques, looking for a setup that is robust and widely applicable. Next, we compare LLMs with skilled human annotators in terms of IAA. Where possible, we contrast the IAA between LLMs and human annotators with the upper bound of the IAA between the annotators themselves. Finally, we manually analyze a subset of LLM and human annotations to assess the validity of the annotations and their respective explanations.
For our experiments, we select three diverse span annotation tasks (cf. Section 3.1): evaluating data-to-text generation Thomson & Reiter (2020), identifying errors in machine translation Kocmi et al. (2024a), and detecting propaganda techniques in human-written texts Da San Martino et al. (2019). Note that unlike established span annotation tasks such as named entity recognition Tjong Kim Sang & De Meulder (2003); Wang et al. (2025a), there are no general purpose fine-tuned models for these tasks since annotation guidelines are specified ad hoc, making their automation difficult.
We base our span annotators on open and proprietary state-of-the-art LLMs, including recent reasoning LLMs that leverage test-time compute for better performance on challenging tasks Muennighoff et al. (2025); Marjanović et al. (2025).
@@
    We demonstrate that zero-shot prompting in combination with structured outputs can be a robust way to set up LLMs as span annotators, given that the model is provided detailed annotation guidelines (Sections 3.2 and 5.1)
@@
    Using a combination of specialized IAA metrics and manual analysis, we show that performance of LLMs is close to skilled human annotators, in some scenarios leveling or surpassing the performance of qualified crowdworkers (Sections 5.2 and 5.3).
@@
    We release a dataset of more than 40k human and model annotations for all our tasks and setups, including reasoning traces from reasoning LLMs and annotations collected from skilled crowdworkers.
@@ 2 Related Work
Automatic NLG metrics traditionally assess text quality by measuring similarity to human-written reference texts Sai et al. (2023); Schmidtová et al. (2024). As such, they are unable to quantify more fine-grained aspects (Gehrmann et al., 2023; Freitag et al., 2021) and do not correlate well with human judgments Novikova et al. (2017); Reiter (2018). With the emerging LLM-as-a-judge paradigm (Gu et al., 2024), LLMs have been applied as evaluators across diverse tasks, either with simple numeric scoring (Bavaresco et al., 2024; Liu et al., 2023; Sottana et al., 2023; Leiter et al., 2023; Chiang & Lee, 2023) or with detailed feedback (Li et al., 2024; Kim et al., 2024a; b; Xu et al., 2023). However, the scores produced by LLMs still miss fine-grained text aspects and are influenced by LLM biases Stureborg et al. (2024); Koo et al. (2024); Wang et al. (2024).
@@
In machine translation (MT), span annotation is a long-standing component of protocols such as MQM or ESA (Lommel et al., 2014; Mariana, 2014; Popovic, 2020; Kocmi et al., 2024c), where annotators mark erroneous spans in translations. In data-to-text (D2T) generation, span annotation is applied by Thomson & Reiter (2020), who introduce a span-based evaluation protocol for annotation of generated basketball reports. Span annotation is also used to judge intrinsic text qualities, such as coherence or use of rhetorical devices, in tasks such as propaganda detection Da San Martino et al. (2019) and text summarization Subbiah et al. (2024). Note that unlike our work, these works perform span annotation only using human annotators.
Automating Span Annotation.
For all the aforementioned tasks, first attempts at automating span annotation were based on fine-tuned pre-trained encoder models. That includes MT (Guerreiro et al., 2024), D2T generation Kasner et al. (2021), text summarization Goyal et al. (2022), and propaganda detection Martino et al. (2020); Goffredo et al. (2023); Piskorski et al. (2023). Automating span annotation with LLMs is more flexible and benefits from increasing LLM capabilities. We build on recent work that applies LLMs as a task-specific evaluation tool (Kocmi & Federmann, 2023; Fernandes et al., 2023; Hasanain et al., 2024; Kasner & Dušek, 2024; Chang et al., 2024; Zouhar et al., 2025; Kartác et al., 2025; Ramponi et al., 2025). Our work is the first that unifies diverse span annotation tasks within a single framework and systematically compares the performance of LLMs to human annotators.
See more discussion on related work in Appendix F.
@@ 3 Automating Span Annotation with LLMs
In this section, we first formally introduce the span annotation process in Section 3.1. Next, we discuss how to automate the process with LLMs in Section 3.2 and how to evaluate the quality of span annotations in Section 3.3.
@@ 3.1 Span Annotation: Task Definition
The goal of span annotation is to annotate a text sequence $Y=\langle y_{1},\ldots,y_{n}\rangle$ given:
@@
  • the annotation guidelines $\mathcal{G}$ (including category definitions and other instructions),
@@
The goal is to produce a set of annotations $A=\{a_{1},\ldots,a_{m}\}$, where each annotation $a_{i}$ is a tuple $\langle s_{i},e_{i},c_{i},r_{i}\rangle$:
@@
  • $r_{i}$ is an explanation for the annotation (optional). [^2] We collect the explanations from LLMs, but not from human annotators due to time considerations.
3.2 Automating Span Annotation with LLMs
To automate span annotation, we collect the annotations $A$ for the given input $\langle Y,C,\mathcal{G},X\rangle$ from an LLM:
$A=\textsc{LLM}(\text{prompt}(Y,C,\mathcal{G},X)).$ (1)
How to Identify the Spans?
An annotation needs to contain the start and end indices of the span. For encoder models, the task would be trivially solved by classifying tokens on top of the encoder’s hidden states. However, to the best of our knowledge, there is no established way of tagging the input text with decoder-only models. Of the methods from previous work, we discarded surrounding spans with special XML-like tokens Wang et al. (2025b); Treviso et al. (2024). This method requires the LLM to generate full copy of the input text and does not offer a fallback if the model fails to do so. We also rule out listing start and end indices of spans Hasanain et al. (2024); Ramponi et al. (2025), as Hasanain et al. note that the model frequently generates indices that do not match the corresponding spans in a paragraph. Instead, we adopted the method of listing textual content of the spans Kocmi & Federmann (2023); Zouhar et al. (2025); Kasner & Dušek (2024). Here, the start and end indices of the span are found heuristically by string matching of the span’s textual content. With this approach, we do not need to modify the input and failing to match a single annotation does not influence other matches. [^3] We do note that this approach can misalign the model output by assigning the annotation to the first match, but our preliminary experiments showed that the collisions are rare.
Parsing Model Output
To obtain the annotations, we request the list of annotations in JSON format, following the setup of Kasner & Dušek (2024). This setup allows us to use constrained decoding with a defined JSON scheme, ensuring that the output is syntactically valid. We require each annotation to contain the fields reason (the explanation $r_{i}$), text (the textual content of the span), and type (the integer index of the error category $c_{i}$). [^4] Given the considerations of Castillo (2024), we ensure that the reason field has precedence.
@@ 3.3 Evaluating Span Annotations
To compare annotations automatically, we need a notion of similarity between two sets of annotations $\mathcal{A}=\{A_{1},A_{2},\ldots,A_{|Y|}\}$ and $\hat{\mathcal{A}}=\{\hat{A}_{1},\hat{A}_{2},\ldots,\hat{A}_{|Y|}\}$ over a set of texts $\mathcal{Y}=\{Y_{1},Y_{2},\ldots,Y_{|Y|}\}$. We consider three similarity metrics:
@@
This metric simply compares how many spans were annotated for each example:
$\mathrm{CountCorrelation}(\mathcal{A},\hat{\mathcal{A}})=\mathrm{Pearson}(\langle|A_{Y}|,|\hat{A}_{Y}|\rangle_{Y\in\mathcal{Y}})$ (2)
This metric serves as a sanity check: a low value would suggest that an annotator either skips examples or over-annotates, indicating unclear annotation guidelines.
@@
To quantify the degree of alignment between individual annotations, we compute precision, recall, and $F_{1}$ as defined in Da San Martino et al. (2019). These measures are on matching annotations, adjusted to give partial credit to imperfect matches: $\text{Precision}(A_{Y},\hat{A}_{Y})=\frac{1}{|A_{Y}|}\sum_{a\in A_{Y}}\frac{|a\cap\hat{a}|}{|a|},$ (3) $\text{Recall}(A_{Y},\hat{A}_{Y})=\frac{1}{|\hat{A}_{Y}|}\sum_{\hat{a}\in\hat{A}_{Y}}\frac{|a\cap\hat{a}|}{|\hat{a}|},$ (4) where $a\cap\hat{a}$ is the character overlap between two annotation spans and $|a|=e-s+1$ is the length of the annotation span in characters (see Section 3.1). Subsequently, we compute the $F_{1}$-score as the harmonic mean of precision and recall.
@@
The $F_{1}$ score is sensitive to varying span granularities and does not consider near-matches with no overlap or agreement by chance. To this end, we follow Da San Martino et al. (2019) and Hasanain et al. (2024) in using the $\gamma$ score (Mathet et al., 2015) as a complementary metric. In short, the metric builds the best possible alignment between the sets of annotations $A_{Y}$ and $\hat{A}_{Y}$ and computes the “disorder” of this alignment based on the positional and categorical dissimilarities of aligned annotations. The score ranges from negative infinity to 1, where 1 is achieved when the annotations are perfectly aligned. The $\gamma$ score extends Krippendorff’s $\alpha$ (Krippendorff, 1980), another popular metric, by computing the category-aware span alignments. We use the implementation of Titeux & Riad (2021).
@@
For an output $y$, one or both annotation sets $A_{Y},\hat{A}_{Y}$ can be empty. Having these cases is desirable: e.g. if the goal is to annotate errors in an output that is entirely correct. However, neither of the introduced scores accounts for these cases. [^5] The $\gamma$ score is undefined for less than two non-empty annotation sets: these examples thus need to be skipped during the $\gamma$ computation. The F1 score does not account for examples with no annotations. To compensate for that, we introduce a score $S_{\emptyset}$ that is computed for examples with less than two nonempty sets of annotations as $S_{\emptyset}=1/(1+|A|)$, where:
@@ 4 Experiments
Table 1: Overview of span annotation tasks used in our experiments. # Cat. denotes the number of categories used in the task (see Appendix C for their listings), # Texts the number of texts annotated, and Avg. Len the average number of words/characters in the output.
@@
We cover three span annotation tasks of different quantitative and qualitative aspects: evaluating data-to-text generation (D2T-Eval; Section 4.1), identifying errors in machine translation (MT-Eval; Section 4.2), and detecting propaganda techniques (Propaganda; Section 4.3). In Table 1, we provide a quantitative overview of our tasks. [^6] See also Figure 1 and Appendix G for the visualization of examples of our tasks.
4.1 Task: Evaluation of Data-to-text Generation
In D2T-Eval, we focus on evaluating the semantic accuracy and stylistic aspects of texts generated from structured data Celikyilmaz et al. (2020); Sharma et al. (2022). The input is the text $Y$ generated from the structured data $X$.
Subtasks and input data.
We select three D2T generation subtasks: generating five-day weather forecasts (openweather), generating product descriptions (gsmarena), and generating reports of soccer games (football). We use the approach of Kasner & Dušek (2024) to download novel structured data from publicly available APIs to make sure that they are unseen for all LLMs (cf. Section 4.4 on data contamination). We provide more details on the data collection process in Section B.1.
Collecting model outputs.
We generate outputs for the structured inputs using LLMs. Specifically, we selected two state-of-the-art models, Llama 3.3 70B Grattafiori et al. (2024) and GPT-4o Hurst et al. (2024), and two smaller models, Gemma 2 2B Team et al. (2024) and Phi 3.5 3.8B Abdin et al. (2024). We prompt the models in a zero-shot setting, asking them to generate a summary of the given data using approximately five sentences. [^7] See Sections A.1 and A.2 for details on the models and Appendix D for the prompts.
Data splits.
For further experiments, we divide the data into multiple subsets:
  • $\mathcal{D}_{\texttt{test}}$ (1200 outputs) – the subset for LLM evaluation, annotated with crowdworkers,
  • $\mathcal{D}_{\texttt{dev}}$ (84 outputs) – the subset for the study of prompt variants, annotated internally,
  • $\mathcal{D}_{\texttt{iaa}}$ (12 outputs) – the control subset for human crowdworkers, annotated internally.
Crowdsourcing annotations

### v3

# LLMs as Span Annotators: A Comparative Study of LLMs and Humans

@@ Abstract
Span annotation – annotating specific text features at the span level – can be used to evaluate texts where single-score metrics fail to provide actionable feedback. Until recently, span annotation was done by human annotators or fine-tuned models. In this paper, we study whether large language models (LLMs) can serve as an alternative to human annotators. We compare the abilities of LLMs to skilled human annotators on three span annotation tasks: evaluating data-to-text generation, identifying translation errors, and detecting propaganda techniques. We show that overall, LLMs have only moderate inter-annotator agreement (IAA) with human annotators. However, we demonstrate that LLMs make errors at a similar rate as skilled crowdworkers. LLMs also produce annotations at a fraction of the cost per output annotation. We release the dataset of over 40k model and human span annotations for further research. [^1] Project website: https://llm-span-annotators.github.io
LLMs as Span Annotators: A Comparative Study of LLMs and Humans
Zdeněk Kasner1   Vilém Zouhar2   Patrícia Schmidtová1 Ivan Kartáč1Kristýna Onderková1Ondřej Plátek1 Dimitra Gkatzia3Saad Mahamood4Ondřej Dušek1Simone Balloccu5 1Charles University 2ETH Zurich 3Edinburgh Napier University 4trivago N.V. 5TU Darmstadt, Germany Contact: kasner@ufal.mff.cuni.cz
@@ 1 Introduction
Fine-grained aspects of texts, such as semantic accuracy or coherence, depend on local lexical choices. To reflect these aspects in quality judgments of texts, techniques are needed that provide the appropriate amount of detail. However, most automatic evaluation metrics for Natural Language Generation (NLG) assign only singular scores for the whole text per each evaluated aspect gkatzia2015snapshot; sai2022survey; schmidtova2024automatic. Although numerical values make it easy to rank systems, these metrics are too crude and susceptible to biases or miscalibration of the underlying models gehrmann2022repairing; liu2023calibrating; wang2023large; gao2024llm.
The subject of our study, span annotation (Figure˜1), offers an alternative approach. Instead of assigning a single score for each evaluated aspect, the goal of span annotation is to localize text spans of interest and classify them according to task-specific guidelines. Span annotations are aligned to specific parts of the evaluated text, which makes them more explainable and actionable than numerical ratings.
Despite its advantages, span annotation has not yet been widely applied in automatic NLG evaluation. The method traditionally required human annotators, making it costly and difficult to scale dasanmartino2019finegrained; thomson2020gold; popovic2020informative; kocmi2024error. The LLM-as-a-judge paradigm recently emerged as a promising solution to this problem zheng2023judging; gu2025surveyllmasajudge, allowing task-specific applications kocmi2023gembamqm; hasanain2024large. However, to our knowledge, no study has systematically compared span annotation performance between LLMs and human annotators.
The central focus of our investigation is comparing human annotators and state-of-the-art LLMs on span annotation tasks. We select three span annotation tasks (cf. Section˜3.1): evaluating data-to-text generation thomson2020gold, identifying errors in machine translation kocmi-etal-2024-findings, and detecting propaganda techniques in human-written texts dasanmartino2019finegrained.
@@
    We establish that with structured outputs and detailed annotation guidelines, LLMs can serve as robust span annotators (Sections˜3.2 and 5.1).
@@
    We show that LLMs have moderate inter-annotator agreement with human annotators overall, but can reach the agreement of qualified crowdworkers (Section˜5.2).
@@
    We discover the sources of model errors by manually analyzing a subset of LLM outputs (Section˜5.3).
  4. 4.
    We release a dataset of more than 40k human and model annotations, including annotations collected from crowdworkers and reasoning traces from reasoning LLMs.
@@ 2 Related Work
Automatic NLG metrics traditionally assess text quality by measuring similarity to human-written reference texts sai2022survey; schmidtova2024automatic. As such, they are unable to quantify more fine-grained aspects (gehrmann2022repairing; 10.1162/tacl_a_00437) and do not correlate well with human judgments novikova-etal-2017-need; reiter-2018-structured. With the emerging LLM-as-a-judge paradigm (gu2025surveyllmasajudge), LLMs have been applied as evaluators on various tasks, using simple numeric scoring (bavaresco2024llms; liu2023geval; sottana2023evaluation; leiter2023eval4nlp; chiang2023can), or free-form feedback (li2024generative; kim2024prometheusa; kim2024prometheus; kartac2025openlgauge). However, as these outputs are not firmly grounded in text, they tend to miss fine-grained aspects and are influenced by LLM biases stureborg2024large; koo2023benchmarking; wang2023large.
@@
In machine translation (MT), span annotation is a long-standing component of protocols such as MQM or ESA (lommel2014multidimensional; mariana2014multidimensional; popovic2020informative; kocmi2024error), where human annotators mark erroneous spans in translations. In data-to-text (D2T) generation, span annotation was applied by thomson2020gold, who introduced a span-based evaluation protocol for annotation of generated basketball match reports. Span annotation is also used to judge intrinsic text qualities, such as coherence or use of rhetorical devices, in tasks such as propaganda detection dasanmartino2019finegrained and text summarization subbiah2024reading. Unlike our work, these works focus on span annotation with human annotators.
Automatic Span Annotation.
Early attempts at automating span annotation with ad-hoc guidelines were based on fine-tuned pre-trained encoder models. That includes evaluation of MT (10.1162/tacl_a_00683), D2T generation kasner2021textincontext or text summarization goyal2022snac, as well as propaganda detection martino2020semeval2020; goffredo2023argumentbased; piskorski2023multilingual. Automating span annotation with LLMs is more flexible and benefits from increasing LLM capabilities. We build on recent work that applies LLMs as a task-specific evaluation tool (kocmi2023gembamqm; fernandes-etal-2023-devil; hasanain2024large; kasner2024traditional; chang2024booookscore; zouhar2024aiassisted; kartac2025openlgauge; ramponi2025finegrained). Our work is the first that unifies diverse span annotation tasks within a single framework and systematically compares the performance of LLMs to human annotators.
@@ 3 Automating Span Annotation with LLMs
We first formally introduce the span annotation process in Section˜3.1. Next, we discuss how to automate the process with LLMs in Section˜3.2 and how to evaluate the quality of span annotations in Section˜3.3.
@@ 3.1 Span Annotation: Task Definition
The aim of span annotation is to annotate a text sequence $Y=\langle y_{1},\ldots,y_{n}\rangle$ given:
@@
  • the annotation guidelines $\mathcal{G}$,
@@
The output is a set of annotations $A=\{a_{1},\ldots,a_{m}\}$, where each annotation $a_{i}$ is a tuple $\langle s_{i},e_{i},c_{i},r_{i}\rangle$:
@@
  • $r_{i}$ is a reason for the annotation (optional).
3.2 Span Annotation with LLMs
In our setup, annotations $A$ for the given input $\langle Y,C,\mathcal{G},X\rangle$ are collected from an LLM:
$A=\textsc{LLM}(\text{prompt}(Y,C,\mathcal{G},X)).$
To obtain the annotations, we follow the setup of kasner2024traditional: we request the list of annotations in JSON format, using constrained decoding with a fixed JSON scheme to ensure that the output is syntactically valid. We require each annotation to contain the fields reason (the explanation $r_{i}$), text (the textual content of the span), and type (the integer index of the error category $c_{i}$). [^2] Following castillo2024structured, we ensure that the reason field is generated first.
For reasoning models not supporting structured output, we retrieve the raw answer from the model, strip any parts within the <think></think> tags (if present), and consider the latest valid top-level JSON object as the model’s response.
@@ 3.3 Evaluating Span Annotations
To compare annotations automatically, we need a notion of similarity between two sets of annotations $\mathcal{A}=\{A_{1},A_{2},\ldots,A_{|Y|}\}$ and $\hat{\mathcal{A}}=\{\hat{A}_{1},\hat{A}_{2},\ldots,\hat{A}_{|Y|}\}$ over a set of texts $\mathcal{Y}=\{Y_{1},Y_{2},\ldots,Y_{|Y|}\}$. We consider the following similarity metrics:
@@
This metric compares how many spans were annotated for each example:
$\operatorname{Pearson}(\mathcal{A},\hat{\mathcal{A}})=\operatorname{\rho}(\langle|A_{Y}|,|\hat{A}_{Y}|\rangle_{Y\in\mathcal{Y}})$ (1)
The correlation serves as a sanity check: a low value would suggest that an annotator either skips examples or over-annotates, indicating unclear annotation guidelines.
@@
To quantify the degree of alignment between individual annotations, we compute precision, recall, and $F_{1}$ as defined by dasanmartino2019finegrained. These measures are on matching annotations, adjusted to give partial credit to imperfect matches:
$\hskip-31.29802pt\operatorname{Precision}(A_{Y},\hat{A}_{Y})=\frac{1}{|A_{Y}|}\sum_{a\in A_{Y}}\frac{|a\cap\hat{a}|}{|a|},\hskip-22.76219pt$ (2)
$\operatorname{Recall}(A_{Y},\hat{A}_{Y})=\frac{1}{|\hat{A}_{Y}|}\sum_{\hat{a}\in\hat{A}_{Y}}\frac{|a\cap\hat{a}|}{|\hat{a}|}$, (3)
where $a\cap\hat{a}$ is the character overlap between two annotation spans and |a|=es+1|a|=e-s+1 is the length of the annotation span in characters (see Section˜3.1). Subsequently, we compute the $F_{1}$-score as the harmonic mean of precision and recall.
@@
The $F_{1}$ score is sensitive to varying span granularities and does not consider near-matches with no overlap or agreement by chance. To this end, we follow dasanmartino2019finegrained and hasanain2024large in using the $\gamma$ score (mathet2015unified) as a complementary metric. The metric builds the best possible alignment between the sets of annotations $A_{Y}$ and $\hat{A}_{Y}$ and computes the “disorder” of this alignment based on the positional and categorical dissimilarities of aligned annotations. The score ranges from -$\infty$ to 1, where 1 is achieved when the annotations are perfectly aligned. The $\gamma$ score extends Krippendorff’s $\alpha$ (krippendorff1980content), another popular metric, by computing the category-aware span alignments. We use the implementation of Titeux2021.
@@
For an output $y$, one or both annotation sets $A_{Y},\hat{A}_{Y}$ may be empty. This is in fact desirable, e.g., if the goal is to annotate errors in an output that is entirely correct. However, these cases are not properly reflected by the other scores we are using: the F1 score only focuses on counting error spans and is not affected by true negatives, and the $\gamma$ score is undefined if any of the two annotation sets is empty (these examples therefore need to be skipped during the $\gamma$ computation). To compensate for this, we introduce a score $S_{\emptyset}$that is computed for examples where any of $A_{Y}$, $\hat{A}_{Y}$ is empty:
$S_{\emptyset}=1/(1+|A|),$ (4)
where:
@@ 4 Experiments
Table 1: Overview of span annotation tasks used in our experiments. # Cat. denotes the number of categories used in the task (see Appendix˜C for their listings), # Texts the number of texts annotated, Avg. Len the average number of words/characters in the output, and Novel Data indicates newly collected data.
4.1 Tasks
@@
We cover three span annotation tasks of different qualitative aspects. We focus on tasks that do not have extensive training data resources and cannot be readily solved by encoder models (such as, e.g., named entity tagging): evaluating data-to-text generation (D2T-Eval; Section˜4.1.1), identifying errors in machine translation (MT-Eval; Section˜4.1.2), and detecting propaganda techniques (Propaganda; Section˜4.1.3). See Table˜1 for an overview of our datasets.
4.1.1 D2T-Eval: Evaluation of Data-to-text Generation
In D2T-Eval, we use span annotation to evaluate semantic accuracy and stylistic aspects of data-to-text generation outputs sharma2022innovations; celikyilmaz2021evaluation. The inputs $X$ are the structured data used to generate the output text $Y$.
We use D2T-Eval as a control task to mitigate the effects of data contamination: the fact that the performance of the model might be inflated by previous exposure to publicly available benchmarks balloccu2024leak; dong2024generalization; jiang2024investigating. Instead of using an existing dataset, we use the Quintd tool kasner2024traditional to download structured inputs from multiple public APIs. [^3] We selected two of the existing domains: openweather for generating weather forecasts and gsmarena for generating phone descriptions. We also add the football domain (using RapidAPI - API-Football) for generating soccer game reports. To obtain output texts for the structured data, we prompt LLMs in a zero-shot setting, asking them to generate a summary of the given data using approximately five sentences. See Appendix B.1 for more details.